OpenStudy (anonymous):

Simplify the expression. Write the result using positive exponents. 3x^3y^3/-8x^2y^9 teach me please?

OpenStudy (anonymous):

x^3 divided by x^2, you get x. y^3 divided by y^9, you get y^-6. since you need to get positive exponents. y^-6 moves to the bottom and becomes y^6. final answer: 3x/8(y^6)

OpenStudy (anonymous):

i forgot the negative sign. sorry! -3x/8(y^6)

OpenStudy (anonymous):

ah..ok

OpenStudy (lalaly):

\[\large{\frac{3x^3y^3}{8x^2y^9}}\] rules of exponents: \[\large{\frac{x^m}{x^n}}= x^{m-n}\]\[\large{x^{-m}=\frac{1}{x^m}}\]

OpenStudy (lalaly):

first factor out the constants \[\large{\frac{3}{8}(\frac{x^3y^3}{x^2y^9}})\]

OpenStudy (lalaly):

now look at the x first since its \[\large{\frac{x^3}{x^2} so =x^{3-2}}=x\]

OpenStudy (lalaly):

then look at the y \[\large{\frac{y^3}{y^9}=y^{3-9} = y^{-6}}\]

OpenStudy (lalaly):

so now you have \[\large{\frac{3}{8}(xy^{-6}})\]

OpenStudy (lalaly):

and u know that \[\large{x^{-m}=\frac{1}{x^m}} so .. y^{-6}=\frac{1}{y^{6}}\]

OpenStudy (lalaly):

there fore the final answer is \[\large{\frac{3x}{8y^6}}\]

OpenStudy (lalaly):

and just put that negative that i forgot :)

OpenStudy (anonymous):

okay thanks